• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 08:03:09 on January23,2003

    Entry number 93086

    keyword=shift summary


    Production ran most of the shift, more or less smoothly. We have almost
    finished 5.2.3, with 96k events of 105k needed.

    Beam position was a bit unstable, with MCC forgetting to restart FFB after
    frequent beam trips.

    6:10 beam gone due to RF failure for about 1 hour.


    -------------------------------------------------------------------------
    20499, 1499 (inherited from swing) 3He production, 5.2.3.
    20500, 1500 3He production, 5.2.3.
    20501, 1501 3He production, 5.2.3.
    20502, 1502 3He production, 5.2.3. Beam position a little unstable
    (MCC kept forgetting to turn on FFB after trips)
    NMR : 40.9%
    20503, 1503 3He production, 5.2.3. starts with bad position
    20504, 1504 3He production, 5.2.3.
    20505, 1505 3He production, 5.2.3.
    -------------------------------------------------------------------------

    00:48 ROC14 crash. reset and re-downloaded L-HRS DAQ.
    01:20 beam gone for tuning (Hall B was having some problem). MCC estimates
    10 minutes.
    01:35 beam returns... it takes a few minutes (and a small reminder from us) for
    MCC to set the position correctly and get FFB back on.
    01:55 beam gone... MCC says that they will need 15-20 minute to correct "some
    beam problems" (I guess they've caught on that I don't know any
    of their jargon...) We ended run 20500, 1500.
    02:11 beam returns
    04:32 beam off for polarization measurement.
    04:41 When beam returns, position measurements are wildly off (4ax=1.7, 4bx=2.5)
    I ask MCC to fix this without high current, and they play with
    beam at 2 uA to re-establish the correct positions.
    05:51 positions are stable, 12 uA on target.... briefly. Very unstable.
    06:12 Informed by MCC that RF cavity problem will cause ~10 minute down.
    06:25 MCC says they need to work around the broken RF cavity. Expect 15 minutes
    down. Run 20504,1504 ended.
    07:06 RF problem fixed, beam returns. Positions are slightly off, but an overtaxed
    operator promises to work on it.